QUESTION 5 Suppose that a researcher, using wage data on 250 randomly selected male workers and 280 female workers, estimates the OLS regression: Wage 12.52 + 2.12 x Male, R2 0.06, SER 4.2 (0.23) (0.36) where Wage is measured in dollars per hour and Male is a binary variable that is equal to 1 if the person is a male and O if the person is a female. Define the wage gender gap as the difference in the mean earnings between men and women. What is the estimated gender gap? (Answer is in $/hour).
QUESTION 6 Suppose that a researcher, using wage data on 250 randomly selected male workers and 280 female workers, estimates the OLS regression: Wage 12.522.12 x Male, R2-0.06, SER 4.2 (0.23) (0.36) where Wage is measured in dollars per hour and Male is a binary variable that is equal to 1 if the person is a male and 0 if the person is a female. Define the wage gender gap as the difference in the mean earnings between men and women. Compute the p-value for testing the null hypothesis that there is no gender gap
QUESTION 7 Suppose that a researcher, using wage data on 250 randomly selected male workers and 280 female workers, estimates the OLS regression: Wage = 12.52 + 2.12xMale, R-0.06, SER = 4.2 (0.23) (0.36) where Wage is measured in dollars per hour and Male is a binary variable that is equal to 1 if the person is a male and O if the person is a female. Define the wage gender gap as the difference in the mean earnings between men and women. If a confidence interval is 1.41 sB1s 2.83, what critical value was used to construct it? O a. 0.67 b. 1.64 OC.1.96 o d. 2.33 e 2.58
QUESTION 8 Suppose that a researcher, using wage data on 250 randomly selected male workers and 280 female workers, estimates the OLS regression: Wage 12.52 2.12xMale, R2 -0.06, SER 4.2 (0.23) (0.36) where Wage is measured in dollars per hour and Male is a binary variable that is equal to 1 if the person is a male and 0 if the person is a female. Define the wage gender gap as the difference in the mean earnings between men and women. In the sample, what is the mean wage of women? (Answer is in S/hour)
Answer #1

5) estimated gender gap =2.12

6) here test statistic =2.12/0.36=5.889

for above test statistic p value =0.0000

7)

as critical value =margin of error/std error =(2.83-1.41)/(2*0.36)=1.972 ~ 1.96

8)

mean wage of women =12.52